# Disentangled Clothed Avatar Generation with Layered Representation

### v1

@@ 3.4 Learning Disentangled Clothed Avatar
Figure 4: Disentanglement and animation capacity. The generated avatars can be decomposed into body, hair, and clothes, which supports expressive animation including gesture and facial expression control.
To mitigate the impact of occlusion, we adopt a single-stage training scheme [6], which is more robust in occluded and sparse view situations. Specifically, the layered UV feature plane fitting and diffusion training process are conducted simultaneously, and the UV feature plane is jointly optimized by the fitting and diffusion loss. Similar to the SDS loss [43], the diffusion loss provides diffusion prior for UV feature plane, thereby facilitating the completion of unseen regions in the training images.
Layered UV Feature Plane Fitting. Given multi-view images, we optimize the layered UV feature plane and shared decoders to reconstruct avatars with disentangle components. The objective function can be divided into reconstruction and constrain part. The reconstruction loss $\mathcal{L}_{\text{recon}}$ can be formulated as follows:
@@
To achieve the disentanglement between exterior components and human body, we not only minimize the color loss $\mathcal{L}_{\text{color}}$ and mask loss $\mathcal{L}_{\text{mask}}$ on the overall rendering result, but also perform supervision on each component. Specifically, we first render the 3D Gaussians corresponding to each component separately to obtain multi-view images of each component. Then, inspired by Clothedreamer [32], instead of blending the rendering results of these components via estimated depth order to obtain the rendering results of clothed avatars, we directly render all the 3D Gaussians to alleviate artifacts caused by the blending process. The silhouette masks of each component and the clothed avatar are obtained similarly. The ground truth of silhouette masks is estimated based on the semantic segmentation results predicted by Sapiens [25]. We apply Huber loss for both $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ following SCARF [17], due to its robustness to the estimated noisy segmentation results. To enhance the details of generated results, we also employ a perceptual loss $\mathcal{L}_{\text{per}}$ to minimize the difference between extracted features of rendered outputs and targeted images.
@@
$\mathcal{L}_{\text{maskin}}=\lambda_{\text{maskin}}\operatorname{ReLU}\left(\mathcal{R}_{m}^{b}(\mathcal{G_{\text{body}}})-\mathcal{M}_{\text{fg}}\right),$ (8)
where $\mathcal{R}_{m}^{b}(\mathcal{G_{\text{body}}})$ represents the rendered silhouette of the human body, and $\mathcal{M}_{\text{fg}}$ is the estimated foreground silhouette mask. When rendering $\mathcal{R}_{m}^{b}(\mathcal{G_{\text{body}}})$, we detach the opacity values and set them to 1, only optimizing offset and covariance-related attributes, preventing the model from minimizing the loss via decreasing opacity on the boundary. Utilizing the prior that the occluded skin color should be similar to the color of hands, we introduce the following texture constrain:
$\mathcal{L}_{\text{skin }}=\lambda_{\text{skin}}\left(\mathcal{M}_{\text{oc}}\odot\left(\mathcal{R}_{m}^{b}(\mathcal{G}_{\text{body}})-\mathrm{C}_{\text{skin}}\right)\right),$ (9)
where $\mathrm{C}_{\text{skin}}$ is the average color of pixels in the hands region and $\mathcal{M}_{\text{oc}}$ is the mask of the occluded region. Other regularization terms are as follows:
@@
$\mathcal{L}_{\text{offset}}=\|\Delta_{\mu}\|^{2}$ constrains the offset from being extremely large. $\mathcal{L}_{\text{smooth}}$ is the total variational(TV) loss, which is used to minimize the average $L_{2}$ distance between neighboring pixels on attribute maps. This regularization term encourages smooth transitions between the neighboring attributes(e.g. offsets, rotation, and opacity), promoting the generation of reasonable geometry surface.
Disentangled Clothed Avatar Diffusion Model. To generate disentangled clothed avatars, we train a diffusion model that maps Gaussian noise to the layered UV latent space. Since diffusion models generally perform better on inputs with low channel dimensions [47], we concatenate our layered UV feature plane across widths instead of stacking them across channels. During training, we first obtain destructed UV feature plane $x_{t}$ by adding Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to the layered UV feature plane $x_{0}$ according to following noise schedule:
$x_{t}:=\alpha(t)x_{0}+\sigma(t)\epsilon,$ (11)
where $\alpha(t)$ and $\sigma(t)$ are predefined functions that control the intensity of added noisy, and $t$ is the time step in the range of [0, 1]. Then, we train a denoising UNet to predict the denoised output via $\hat{x}_{0}=\alpha(t)x_{t}-\sigma(t)\hat{v}$, according to the $v$ parameterization technique in  [48]. The subjective function for the diffusion model can be expressed as:
$\mathcal{L}_{\text{diff}}=\underset{t,x_{0},\epsilon}{\mathbb{E}}\left[\frac{1}{2}w(t)\left\|\ \hat{x}_{0}-x_{0}\right\|^{2}\right],$ (12)
where $w(t)=\left(\alpha(t)/\sigma(t)\right)^{2\omega}$, the $\omega$ is set to 0.5 in practice. $\mathcal{L}_{\text{diff}}$ is used to update not only the parameters of denoising UNet, but also the UV feature plane. It promotes the UV feature plane to adapt to the learned latent space, thereby providing priors for occluded regions.
@@ 4 Experiments
Figure 5: Qualitative comparison on Tightcap [9] and THuman2.0 [66] Dataset. The left part illustrates layer-wise generated results on Tightcap dataset [9]. Our method can generate more reasonable human bodies under severe occluded scenarios. The right part demonstrates our generation results on THuman2.0 dataset [66]. Our method generate high-quality results with clear fingers and cloth wrinkles.
Dataset. We evaluate our method on two datasets: Tightcap [9], and THuman2.0 [66]. Tightcap contains 3D scans with separate meshes for cloth and shoe components, providing easy extraction to multi-view rendering images and silhouette masks for each component. We use the preprocessed version provided by HumanLiff [21], which selects 107 samples for Tightcap. Each sample is rendered from 158 camera views using the Blender engine, providing RGB images and silhouette masks for the entire clothed avatar and each component. To evaluate the capacity of our model to learn from multi-view images without the need for separated meshes for each component, we further conduct experiments on the THuman2.0 dataset. Specifically, we sample 500 scans and render from 54 camera views to obtain the multi-view images. Then, we employ Sapiens [25] to estimate segmentation masks for these images, where silhouette masks for each component can be extracted.
4.1 Evaluation on Generated Clothed Avatar
Disentanglement and Animation Capacities. Our method can generate clothed avatars with full disentanglement of components such as hair, shoes, and clothes with inner human body. The disentanglement and animation results are shown in Fig 4, our method can separate each component and learn the inner body layer with reasonable texture and geometry even under severe occluded scenarios. Benefiting from the SMPL-X-based templates within the layered Gaussian-based UV feature plane representation, each component can be easily deformed into novel poses. Facial expressions and gestures controllability is also supported, demonstrated by exaggerated expressions and distinct finger animation, as shown in Fig 4.
Comparisons. We compare our method with current representative feed-forward diffusion pipelines and well-known 3D-aware GAN-based generation pipelines. The quantitative comparison results are shown in Table 1. Our method outperforms other methods on both TightCap [9] and THuman2.0 [66] dataset, while achieving disentanglement capabilities. The visual comparison is shown in Fig. 5. We evaluate the layer-wise generation results on the Tightcap dataset and the overall generation quality on the THuman2.0 dataset. For layer-wise generation, our method can generate correct results for each layer thanks to the complete disentanglement of each component. EG3D [5], EVA3D [19], and Rodin [59] fail to generate proper components for each layer, and identity shift problem occurs during the process. HumanLiff [21] achieves correct layer-wise generation, but the human body layer is less smooth and reasonable. Compared with other methods on the THuman2.0 dataset, our method can generate clothed avatars with intricate details such as cloth wrinkles and fingers. We further achieve expressive animations, including gesture and expression control and component disentanglement.
Methods TightCap [9] THuman2.0 [66]
@@ 4.2 Ablation Study
Table 2: Ablation study on THuman2.0 Dataset. The full pipeline outperform the baselines on both FID and KID for a large margin.
Figure 6: Ablation study on THuman2.0 Dataset. Comparing randomly generated avatars and decomposition results, our method generates higher quality avatars with better disentanglement.
Layered UV Feature plane. To evaluate the effectiveness of the layered UV feature plane, we design two single-layer UV feature plane representations to generate disentangled clothed avatars. 3D Gaussians belonging to each component are initialized on the SMPL-X surface, thus 3D Gaussians belonging to different components share one feature if they are initialized on the same location. One representation is SLMO(single-layer-multi-output), in which we utilize the geometry and texture decoder $\mathcal{D}_{g}$ and $\mathcal{D}_{t}$ to predict the attribute value of 3D Gaussians belonging to different components with one single feature extracted from the UV feature plane. Another one is SLMN(single-layer multi-network), where for one feature, we utilize different decoders to predict the attribute values of Gaussian primitives belonging to different components. As shown in Fig 6 and Tab 2, single-layer UV representations generate clothed avatars with lower quality both quantitatively and qualitatively. Without the layered UV templates which provide different features for components in neighboring regions, the components are difficult to disentangle, which leads to blurry results and artifacts.
Single Stage Diffusion. We also compare the single-stage training scheme with the commonly used two-stage training scheme. The single-stage training can generate avatars with finer details, such as clearer faces and cloth wrinkles.
@@ 4.3 Applications
Figure 7: Component transfer. Given the generated avatars in the first column, we can transfer the upper-clothes, pants, hair, shoes of the avatars in the second to fifth column to them. The results are shown in the rightmost column.
Component Transfer. We further explore the applications such as component transfer. Thanks to the component disentanglement and shared structure provided by our method, we can directly transfer clothes and other components between generated samples. The component transfer results are shown in Fig 7. Our method can accurately transfer components between subjects in various body shapes while maintaining high-quality results with intricate details.
@@ 5 Conclusion
In this paper, we propose LayerAvatar, the first feed-forward diffusion-based method to generate component-disentangled clothed avatars. Specifically, we first propose a layered UV feature plane representation, which organizes the 3D Gaussians and represents each component of the clothed avatars in different layers with corresponding semantic labels. Then, we train a single-stage diffusion model to generate each feature plane to obtain clothed avatars with each component disentangled. Semantic-aware compositional rendering and constraint terms are elaborately designed for the full disentanglement of each component. Extensive experiments demonstrate the superiority of LayerAvatar in fully disentangled clothed avatar generation and component transfer.

### v2

@@ 3.4 Learning Disentangled Clothed Avatar
To mitigate the impact of occlusion, we adopt a single-stage training scheme [6], which is more robust in occluded and sparse view situations. Specifically, the layered UV feature plane fitting and diffusion training process is conducted simultaneously, and the UV feature plane is jointly optimized by the fitting and diffusion loss. Similar to the SDS loss [55], the diffusion loss provides a diffusion prior for the UV feature plane, thereby facilitating the completion of unseen regions in the training images.
Layered UV Feature Plane Fitting. Given multi-view images, we optimize the layered UV feature plane and shared decoders to reconstruct avatars with disentangled components. The objective function can be divided into reconstruction and constraint part. The reconstruction loss $\mathcal{L}_{\text{recon}}$ can be formulated as follows:
@@
To achieve the disentanglement between exterior components and human body, we not only minimize the color loss $\mathcal{L}_{\text{color}}$ and mask loss $\mathcal{L}_{\text{mask}}$ on the overall rendering result, but also perform supervision on each component. Specifically, we first render the 3D Gaussians corresponding to each component separately to obtain multi-view images of each component. Then, inspired by Clothedreamer [41], instead of blending the rendering results of these components via estimated depth order to obtain the rendering results of clothed avatars, we directly render all the 3D Gaussians to alleviate artifacts caused by the blending process. The silhouette masks of each component and the clothed avatar are obtained similarly. The ground truth of silhouette masks is estimated based on the semantic segmentation results predicted by Sapiens [31]. We apply Huber loss [27] for both $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ following SCARF [18], due to its robustness to the estimated noisy segmentation results. To enhance the details of generated results, we also employ a perceptual loss $\mathcal{L}_{\text{per}}$ [29] to minimize the difference between extracted features of rendered outputs and targeted images.
@@
$\mathcal{L}_{\text{maskin}}=\lambda_{\text{maskin}}\text{ReLU}\left(\mathcal{R}_{m}^{b}(\mathcal{G_{\text{body}}})-M_{\text{fg}}\right),$ (8)
where $\mathcal{R}_{m}^{b}(\mathcal{G_{\text{body}}})$ represents the rendered silhouette of the human body, and $M_{\text{fg}}$ is the estimated foreground silhouette mask. When rendering $\mathcal{R}_{m}^{b}(\mathcal{G_{\text{body}}})$, we detach the opacity values and set them to 1, only optimizing offset and covariance-related attributes, preventing the model from minimizing the loss via decreasing opacity on the boundary. Utilizing the prior that the occluded skin color should be similar to the color of hands, we introduce the following texture constraint:
[…]
where $\mathbf{C}_{\text{skin}}$ is the average color of pixels in the hands region and $M_{\text{oc}}$ is the mask of the occluded region. Other regularization terms are as follows:
@@
$\mathcal{L}_{\text{offset}}=\|\Delta_{\mu}\|^{2}$ constrains the offset from being extremely large. $\mathcal{L}_{\text{smooth}}$ is the total variational (TV) loss, which is used to minimize the average $L_{2}$ distance between neighboring pixels on attribute maps. This regularization term encourages smooth transitions between the neighboring attributes (e.g. offsets, rotation, and opacity), promoting the generation of reasonable texture and geometry surface.
Disentangled Clothed Avatar Diffusion Model. To generate disentangled clothed avatars, we train a diffusion model that maps Gaussian noise to the layered UV latent space. Since diffusion models generally perform better on inputs with low channel dimensions [59], we concatenate our layered UV feature plane across widths instead of stacking them across channels. During training, we first obtain destructed UV feature plane $\boldsymbol{x}_{t}$ by adding Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ to the layered UV feature plane $\boldsymbol{x}_{0}$ according to following noise schedule:
$\boldsymbol{x}_{t}:=\alpha(t)\boldsymbol{x}_{0}+\sigma(t)\boldsymbol{\epsilon},$ (11)
where $\alpha(t)$ and $\sigma(t)$ are predefined functions that control the intensity of added noise, and $t$ is the time step in the range of [0, 1]. To stabilize and accelerate training, we use v-prediction proposed in  [61] to train the denoising UNet. The objective function for the diffusion model is:
$\mathcal{L}_{\text{diff}}=\underset{t,\boldsymbol{x}_{0},\boldsymbol{\epsilon}}{\mathbb{E}}\left[\frac{1}{2}w(t)\left\|\ \hat{\boldsymbol{x}}_{0}-\boldsymbol{x}_{0}\right\|^{2}\right],$ (12)
where $w(t)=\left(\alpha(t)/\sigma(t)\right)^{2\omega}$, the $\omega$ is set to 0.5 following  [6]. $\mathcal{L}_{\text{diff}}$ is used to update not only the parameters of denoising UNet, but also the UV feature plane. It promotes the UV feature plane to adapt to the learned latent space, thereby providing priors for occluded regions.
@@ 4 Experiments
Ours 12.50
[…]
EVA3D [22] 61.58* <<20*
Rodin [72] 56.57* 18.12*
HumanLiff [24] 54.39* 28.57*
Ours 17.37 >>40
[…]
SO-SMPL [70] 13.02 3.57/8.57 5h
LAGA [19] 9.77 5.00/1.43 1.5h
TELA [15] 19.07 14.29/8.57 6h
Ours 58.14 77.14/81.43 2s
Table 1: Quantitative comparison. Our method outperforms other methods in both holistic and layer-wise generation quality, as well as disentanglement capability. Due to the minor difference between the two masked layers, the L-PSNR value of our method is so large that we use >> 40 to express it. For component comparison, we show the quality preference in component quality preference/overall quality preference format. *, $\dagger$} and $\star$ denote results adopted from HumanLiff [24], StructLDM [26] and E3Gen [87] respectively.
Figure 3: Qualitative comparison. The left part demonstrates our generation results on THuman2.0 dataset [81]. The middle part illustrates layer-wise generated results on Tightcap dataset [9]. Our method can generate more reasonable human bodies under severe occluded scenarios. The right part exhibits our component generation results compared with optimization-based methods.
Baselines. To evaluate holistic generation quality, we compare our method with the state-of-the-art methods of animatable avatar generation (EVA3D [22], StructLDM [26], and E3Gen [87]) on THuman2.0 [81] dataset. We then evaluate the decomposition capability of our method by comparing the layer-wise generation results with the disentanglement-related method HumanLiff [24] on Tightcap [9] dataset. And we further evaluate the component generation quality against optimization-based methods: LAGA [19], SO-SMPL [70], and TELA [15]. Our method is trained on a composite dataset including CustomHuman [21], THuman2.0 [81], and THuman2.1 [81].
Metrics. For holistic generation quality evaluation, we utilize FID [20] following previous works [5, 26, 87]. For layer-wise generation results, we adopt FID for overall generation quality evaluation and L-PSNR [24] for disentanglement capability evaluation, which calculates the PSNR between two layers with the adding component masked. Additionally, we conduct a user study with 20 participants to compare the generation quality and disentanglement quality of our methods with optimization-based methods.
Dataset. For THuman2.0 [81], we sample 500 scans and render each from 54 camera views to obtain multi-view images. Then, we employ Sapiens [31] to estimate segmentation masks for these images, from which per-component silhouette masks can be extracted. This approach enables our method to learn directly from multi-view images without requiring separate meshes of each component, thereby simplifying the data collection process. Tightcap [9] contains 3D scans with separate meshes for cloth and shoe, which facilitates direct rendering of multi-view images and silhouette masks for each component. For a fair comparison, we use the preprocessed version provided by HumanLiff [24], which selects 107 samples from Tightcap. Each sample is rendered from 158 camera views, providing images and silhouette masks for both the entire avatar and each component. We also construct a composite dataset consisting of 1954 selected scans from THuman2.0 [81], THuman2.1 [81], and CustomHuman [21] datasets. Each scan is processed in the same way as the THuman2.0 data. Additionally, all SMPL-X parametric models are standardized to neutral gender with refinement in body shape parameters to ensure the body layer fits underneath the cloth surface.
4.1 Evaluation of Generation Quality and Disentanglement Capability
Disentanglement and Animation Capacities. The disentanglement and animation results are shown in Fig. 1. Our method successfully generates clothed avatars with full disentanglement of components such as hair, shoes, and clothes from the human body. Moreover, our method reconstructs the inner body layer with reasonable geometry and texture under severe occlusion. Benefiting from the SMPL-X-based templates within our layered Gaussian-based UV feature plane representation, each component can be easily deformed into novel poses.
Comparisons. For holistic generation quality, we compare our method against representative feed-forward diffusion and 3D-aware GAN pipelines. The quantitative comparison results are shown in Tab. LABEL:tab:full_generation_quality. Our method outperforms all baselines on THuman2.0 [81] dataset, achieving the best FID score. The visual comparisons shown in Fig. 3 further enhance the superiority of our method. EVA3D [22] struggles with generating digital avatars with plausible texture and geometry. Compared to StructLDM [26] and E3Gen [87], our method exhibits finer details, such as distinctive fingers, realistic cloth wrinkles, and intricate faces.
To evaluate disentanglement capability, we compare layer-wise generation results and disentangled avatar generation results of our method with state-of-the-art methods. For layer-wise generation, our method surpasses other methods in both FID and L-PSNR, as shown in Tab. LABEL:tab:layer_generation_quality, which indicates that our method achieves higher holistic generation results and can generate disentangled avatars without facing the identity shifting problem. The elimination of identity shifting demonstrates that our method achieves full disentanglement of components, since entangled components will result in shifting problems. As shown in Fig. 3, Rodin [72] struggles to maintain identity consistency during the generation process. Although HumanLiff [24] achieves correct layer-wise generation, it produces less smooth body geometry and exhibits subtle identity shifting, shown by the enlarged face region. Our method can generate each layer without shifting problems, demonstrating more complete disentanglement.
For disentangled avatar generation, we compare our method with state-of-the-art optimization-based methods. As shown in Tab. LABEL:tab:compo_generation_quality, our method generates clothed avatars in significantly less time and achieves a higher user preference. Qualitative comparisons are illustrated in Fig. 3. SO-SMPL [70] generates unrealistic, cartoonish colors and exhibits unnatural sawtooth on the border of generated components (highlighted by orange circles). LAGA [19] and TELA [15] struggle with blurry fingers, oversaturated colors, and incomplete disentanglement from the body layer. In contrast, our method generates fully disentangled avatars with distinctive fingers and realistic color.
@@ 4.2 Ablation Study
Table 2: Ablation study on THuman2.0 Dataset. The full pipeline outperforms baselines on both FID and KID by a large margin.
Figure 4: Ablation study on THuman2.0 Dataset. Comparing randomly generated avatars and decomposition results, our method generates avatars with higher quality and better disentanglement.
To demonstrate the effectiveness of our layered UV feature plane, we design two baselines that use single-layer UV feature plane to generate disentangled clothed avatars. One baseline representation is SLMO (single-layer-multi-output), which utilizes a single pair of geometry and texture decoders ($\mathcal{D}_{g}$ and $\mathcal{D}_{t}$) to predict attributes for all components. Another one is SLMN (single-layer multi-network), which employs different decoders to predict the attributes of Gaussian primitives for each component. Since they only contain a single-layer UV feature plane, attributes of different components will share one feature if they are initialized in the same location. As shown in Fig. 4, Fig.5 and Tab. 2, single-layer UV representations generate clothed avatars with lower quality both quantitatively and qualitatively. This is because our layered UV feature plane representation provides each component with an optimized UV distribution and independent feature space. Without the layered UV plane, which provides distinctive features for components in overlapping and neighboring regions, single-layer representations struggle to disentangle components, leading to blurred boundaries between components.
Figure 5: Zoom in comparison between E3Gen and LayerAvatar. Compared to single-layer representation, layered representation exhibits detailed faces and clear boundaries between components.
Single Stage vs. Two Stage. We also compare the single-stage training scheme with the commonly used two-stage training scheme, shown in Tab. 2 and Fig. 4. The single-stage training can generate avatars with finer details, such as intricate faces.
@@ 4.3 Applications
Component Transfer.
Figure 6: Component transfer. Given the generated avatars in the first column, we can transfer the upper-clothes, pants, hair, and shoes of the avatars in the second to fifth column to them. The results are shown in the rightmost column.
We further explore applications such as component transfer. Thanks to the component disentanglement and shared structure provided by our method, we can directly transfer clothes and other components between generated samples. The component transfer results are shown in Fig. 6. Our method can accurately transfer components across various body shapes while maintaining high-quality details of the transferred items.
@@ 5 Conclusion
In this paper, we propose LayerAvatar, a novel feed-forward diffusion-based method for generating component-disentangled clothed avatars. Our method proposes a layered UV feature plane representation, which organizes 3D Gaussians into different layers, each corresponding to specific components of clothed avatars (e.g., body, hair, clothing). Leveraging this representation, we train a single-stage diffusion model to generate each feature plane, enabling the generation of fully disentangled clothed avatars. To ensure complete component disentanglement, we incorporate constraint terms into the model. Extensive experiments validate the superiority of LayerAvatar in generating fully disentangled clothed avatars and its effectiveness in component transfer tasks. Additional limitations and implementation details are discussed in the supplementary material.
Acknowledgments.  This work was supported in part by NSFC (62201342), and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).
Supplementary Material
This is the supplementary material for Disentangled Clothed Avatar Generation via Layered Representation. A video (Sec 6) is included summarizing our method and exhibiting more visualization results. We introduce the implementation details of our method in Sec 7. Additional experimental results and limitation discussions are provided in Sec 8.
6 Supplementary Video
We provide a supplementary video for quick understanding of our method. The video includes:
  • A brief introduction of our method;
  • Results of unconditional generation and decomposition;
  • Results of novel pose animation;
  • Results of component transfer.
7 Implementation Details
7.1 Network Architecture
Layered UV Feature Plane and Shared Decoders. The size of layered UV feature plane is $12\times 128\times 384$, where we concatenate the three-layer Gaussian-based UV feature plane width-wise instead of channel-wise following  [72]. Two shared decoders $\mathcal{D}_{g}$ and $\mathcal{D}_{t}$ are utilized to decode the UV feature plane to attribute maps, where attribute (position $\mu$, opacity $\alpha$, rotation $\mathbf{r}$, scale $\mathbf{s}$, color $\mathbf{c}$) of 3D Gaussians [30] can be extracted via bilinear interpolation. $\mathcal{D}_{g}$ predict geometry-related attributes (position and opacity) while $\mathcal{D}_{t}$ outputs texture-related attributes (color, rotation and scale). The architecture of these two shared decoders is shown in Fig. G. $\mathcal{D}_{g}$ and $\mathcal{D}_{t}$ both are shallow decoder with two layers. For the first layer, we apply SiLU [17] as the activation function, while for the last layer, Sigmoid is utilized except for the offset $\Delta\mu$ prediction layer. No activation function is utilized for the offset prediction layer. Both the offset prediction layer and covariance prediction layer (predict $\Delta\mathbf{r}$ and $\Delta\mathbf{s}$) are initialized with weight conform to $\mathcal{U}(-1\times 10^{-5},+1\times 10^{-1})$. Biases are initialized to be 0.
Figure G: Network Architecture of Shared Decoder. The number above the arrow represents the input and output channels of each block. Each block consists of one 2D convolutional layer, one batchnorm layer, and one activation function. The number on the block is the kernel size of the convolutional layer. We utilize SiLU as the activation function and for the last layer, sigmoid is used as the activation function except for the $\Delta\mu$ output.
Denoising UNet. Following [58], the denoising UNet [60] has UNet architecture with attention modules. The network architecture is shown in Tab. C.
Figure H: Structure of Single Layer representations. We design two kinds of single-layer representations for ablation study.
Table C: Model Configuration of denoising UNet.
Key Value
Number of timesteps 1000
Noise configuration Linear
Input size $128\times 384\times 12$
Base channel 128
Channels configuration [0.5, 1, 2, 2, 4, 4]
Resblocks per downsample 2
dropout 0
Number of attention heads 4
Attention resolution [16, 8, 4]
7.2 Training Details
The whole training framework is built upon Pytorch. We utilize Adam [33] as the optimizer. For shared decoder and denoising UNet, the learning rate is $1\times 10^{-4}$ and $1\times 10^{-4}$ separately. The learning rate for UV feature plane is $0.04$. All the UV feature planes are first normalized through a Tanh function and then sent into the shared decoder. The total batch size is 4 scenes per GPU. During one iteration, we select 2 views of each scene for supervision. The whole training process takes 6 days on two RTX 3090 GPUs. The loss weight for each fitting loss is as follows: $\lambda_{\text{color}}=18$, $\lambda_{\text{mask}}=9$, $\lambda_{\text{per}}=0.05$, $\lambda_{\text{seg}}=9$, $\lambda_{\text{maskin}}=5$, $\lambda_{\text{skin}}=0.5$, $\lambda_{\text{offset}}=5$, and $\lambda_{\text{smooth}}=0.5$. For ablation study, we propose single-layer representations with a size of $128\times 128\times 12$, the whole architecture is shown in Fig. H. Other parameters are the same as layered UV feature plane.
7.3 Data Preprocessing
For Tightcap [9], we convert the estimated SMPL [43] parameters to SMPLX [52] to fit our template. Masks for each component (top, bottom, shoes) are provided by the dataset. Since the dataset does not separate hair from body, we combine hair and body components during training. For THuman2.0, THuman2.1 [81] and CustomHuman [21], we first optimize the SMPL-X parameters to make them underneath the clothing layer. We then obtain the segmentation maps for each view through Sapiens [31] estimation and merge the original 20 segmentation labels into 5 (hair, body, top, bottom, and shoes).
8 Limitations and Discussions
8.1 Limitations
(1) Due to the segmentation-map-based supervision and SMPLX-based templates, the performance of our method is affected by the accuracy of the estimated segmentation map and SMPLX parameters. Eliminating inaccurate segmentation results as in  [28] and optimizing SMPLX parameters during training can help mitigate the impact.
(2) The collision between body and clothing is a long-existing problem when representing the human body and clothing separately. Introducing post-processing to optimize the position of 3G Gaussians on clothing and body via collision-avoiding loss might eliminate this problem.
(3) Animation of loose clothing is prone to artifacts. Using video data instead of multi-view images or introducing pre-trained networks for cloth deformation might help.
(4) While our method effectively disentangles core components, a promising direction for future research is to extend it to handle general accessories (e.g., glasses, bags) by incorporating additional template layers or adopting other representations.
(5) Introducing physical constraints by converting current representation to meshes [35] or combining UV feature plane with sewing pattern similar to GarmageNet [36] would be interesting future work.
8.2 Additional Experiments
Shape Adaptation. Benefiting from the SMPL-X-based template of our representation, we can adapt clothing to various body shapes as shown in Fig. I which facilitates seamless transfer of components among subjects.
Figure I: Clothing Adaptation to Body Shape
Sensitivity on noisy semantic masks. Our method is robust to noisy masks thanks to multi-view supervision, which can correct errors present in individual views. During training, we filter out samples with obvious noise in over three consistent views. As shown in Fig. J, our method can handle noise that does not persist across multiple views.
Figure J: Robustness to noisy semantic masks.
Extension to multi-layer outfits. As shown in Fig. K, our method can composite multi-layer clothes to enable more composition.
Figure K: Multi-layer garment composition.
More Generation Results. By training on the composition of three datasets (THuman2.0, THuman2.1, and CustomHuman), our method is capable of generating more diverse avatars and challenging clothing which demonstrates the promise of our methods to achieve better results on larger datasets, as illustrated in Fig. L and Fig. M.
Figure L: More diverse generation results.
Figure M: Dress and skirt generation results.
Generalization capability. As shown in Fig N, our method can generalize to various clothing types (dress) and poses (hitting) exhibiting distinctive fingers, facial details, and disentanglement capability. We believe that better results can be obtained when extending to a larger training set.
Figure N: Single image reconstruction on unseen 4DDress dataset.
Ablation on TVLoss. Shown in Fig. L, without TVLoss, the inner body color will be affected by outer components.
Figure O: Ablation on feature map resolution and decoder depth.
Ablation on UV feature plane resolution and decoder depth. We conduct a simple experiment to explore the influence of UV feature plane resolution and decoder depth on the results. We reconstruct digital avatars utilizing multi-view images with different resolutions and decoder depths. Demonstrated by Fig. O, the increase of feature map resolution can enhance the final results by a small margin. Increasing decoder depth, however, hinders optimization and reduces results. We consider imbalanced data distribution(over 90% plain texture clothing) and occlusion from limited poses might be the primary constraint of the generation quality for our method. Incorporating video and synthetic data might be helpful to construct a more powerful model.